# A Causal Perspective on Measuring, Explaining and Mitigating Smells in LLM-Generated Code

### v2

@@ 1. Introduction
LLMs have rapidly moved from experimental tools to everyday assistants in software engineering (SE) (watson2021systematicliteraturereviewuse). Developers rely on them for many tasks such as code completion (9616462; lin2024soen101codegenerationemulating; 7180092), summarization (ahmed2022fewshottrainingllmsprojectspecific), program repair (jin2023inferfixendtoendprogramrepair), clone detection (10.1145/2970276.2970326) and test generation (Watson_2020). As generated code increasingly enters production systems, evaluation must look beyond functional correctness and consider whether the output meets broader standards of maintainable and robust software.
Code smells offer a foundation for evaluating these broader quality concerns. They capture recurring design and implementation issues that affect readability, complexity, and long-term maintainability (smells_michele; palomba2018diffuseness; Evtikhiev_2023). Prior work has traditionally attributed the introduction of smells to human-driven factors such as developer experience, time pressure, and project constraints (smells_michele). The increasing involvement of LLMs challenges this assumption. Models trained on large-scale code corpora inherit patterns present in their training data, including smelly or suboptimal code (10795572; tambon2024bugslargelanguagemodels), and they may introduce additional defects or vulnerabilities during generation (10.1145/3605098.3636058; pearce2021asleepkeyboardassessingsecurity). Smell formation now reflects both human design choices and model behavior, which makes traditional explanations incomplete.
Current evaluation practices are not designed to capture this shift. Similarity metrics such as BLEU (papineni2002bleu) and CodeBLEU (codebleu), along with execution-based measures, offer useful information about functional correctness but provide limited insight into structural or design quality (empirical_code_smells; 10555682; nunes2025evaluatingeffectivenessllmsfixing; takerngsaksiri2025codereadabilityagelarge). Although static analysis tools (SATs) can identify smells in model outputs, post-hoc detection alone cannot reveal why a model introduced these issues or which components of the generation process shaped them. As a result, developers lack visibility into the factors that influence the maintainability of LLM-generated code.
@@
While these questions remain open, recent work has begun to introduce early indicators of smell propensity in LLMs. Velasco et al. proposed the PSC metric, a probabilistic score that estimates the likelihood that a model emits tokens associated with a given smell (velasco2025propense). Introduced as part of a benchmarking effort, PSC provides a continuous signal of model tendencies and a promising foundation for investigating the factors behind smell occurrence. Yet the broader dynamics of smell generation remain unexplored. The drivers of variation in PSC and their relationship to the structural quality of generated code are still unknown, and the stability of the metric across diverse conditions remains untested. In addition, PSC has yet to be used to uncover mechanisms that explain why models introduce smells or to support interventions that reduce smell propensity. These gaps hinder a full understanding of how smells emerge in LLM-generated code and limit our ability to devise approaches for reducing their occurrence.
Responding to these gaps, we investigate smell generation from a causal perspective. Instead of treating PSC as a standalone benchmark metric, we use it as an instrument for reasoning about the factors that influence smell propensity in LLM-generated code. Our study begins by assessing the robustness and explanatory value of PSC, establishing that it provides a stable foundation for causal analysis. We then quantify the effects of key elements in the generation process, including decoding strategy, model size, model architecture and prompt formulation, on the likelihood of smell occurrence. These causal estimates identify the factors that meaningfully influence smell propensity and guide the design of mitigation strategies that can be applied during inference. Finally, we complement this analysis with a user study that evaluates whether PSC helps developers interpret and assess generated code.
@@
We provide a causal analysis quantifying the influence of generation strategy, model size, model architecture and prompt formulation.
@@
We release a dataset, benchmark and implementation to support future research on explaining and mitigating smell generation (psc_repo).
@@ 4. Methodology
[Usefulness] How do developers assess code smells in LLM-generated code when supported by PSC? We conduct a user study to examine whether PSC helps developers evaluate generated code, focusing on its perceived relevance, its impact on confidence and its usefulness in identifying introduced smells.
@@ 4.1. Semantic Preserving Code Transformations
Our methodology draws on prior work in robustness evaluation for neural program repair using natural transformations (sept_paper), but we focus specifically on statement-level transformations. Broader edits, such as converting while loops to for loops (While2For), can alter structural context and influence how smells are detected. Limiting the analysis to statement-level changes preserves semantics and overall structure while still introducing meaningful syntactic variation.

### v3

@@ 1. Introduction
LLMs have rapidly moved from experimental tools to everyday assistants in software engineering (SE) (watson2021systematicliteraturereviewuse). Developers rely on LLM-based tools for many tasks such as code completion (9616462; lin2024soen101codegenerationemulating; 7180092; sun2025bridgingdeveloperinstructionscode), summarization (ahmed2022fewshottrainingllmsprojectspecific; 11029737; 10.1145/3728949), program repair (jin2023inferfixendtoendprogramrepair; 10.1145/3650212.3680323), clone detection (10.1145/2970276.2970326; 10556419), and test generation (Watson_2020; 10329992; 10765033). As generated code increasingly enters production systems, evaluation must look beyond functional correctness and consider whether the output meets broader standards of maintainable and robust software.
Code smells provide a foundation for evaluating broader quality concerns by capturing recurring design and implementation issues that affect readability, complexity, and long-term maintainability (smells_michele; palomba2018diffuseness; Evtikhiev_2023). Prior work has traditionally attributed smell introduction to human-driven factors such as developer experience, time pressure, and project constraints (smells_michele). The growing involvement of LLMs challenges this assumption. Models trained on large-scale code corpora inherit patterns present in the training data, including smelly or suboptimal code (10795572; tambon2024bugslargelanguagemodels), and may introduce additional defects or vulnerabilities during generation (10.1145/3605098.3636058; pearce2021asleepkeyboardassessingsecurity). Smell formation now reflects a combination of human design choices and model behavior, which renders traditional explanations incomplete.
Current evaluation practices are not designed to reflect model decisions influencing smell generation. For instance, similarity metrics such as BLEU (papineni2002bleu) and CodeBLEU (codebleu), along with execution-based measures, offer useful information about functional correctness but provide limited insight into structural or design quality (empirical_code_smells; 10555682; nunes2025evaluatingeffectivenessllmsfixing; takerngsaksiri2025codereadabilityagelarge). Although static analysis tools (SATs) can identify smells in model outputs, post-hoc detection alone falls short in revealing why a model introduced smells or which components of the generation process shaped them. As a result, developers lack visibility into factors that influence the maintainability of LLM-generated code.
@@
While these questions remain open, recent work has begun to introduce early indicators of smell propensity in LLMs. Velasco et al. proposed the PSC metric, a probabilistic score that estimates the likelihood that a model generates tokens associated with a given smell (velasco2025propense). Introduced as part of a benchmarking effort, PSC provides a continuous signal of model tendencies and a promising foundation for investigating the factors behind smell occurrence. Yet the broader dynamics of smell generation remain unexplored. 1) The drivers of variation in PSC and their relationship to the structural quality of generated code are still unknown, and 2) the stability of the metric across diverse conditions remains untested. Finally, 3) PSC has yet to be used to uncover mechanisms that explain why models introduce smells or to support interventions that reduce smell propensity. These three gaps hinder a full understanding of how smells emerge in LLM-generated code and limit our ability to devise approaches for reducing their occurrence.
Responding to these gaps, we investigate smell generation from a causal perspective. Instead of treating PSC as a standalone benchmark metric, we use it as an instrument for reasoning about the factors that influence smell propensity in LLM-generated code. Our study begins by assessing the robustness and explanatory value of PSC, establishing that it provides a stable foundation for causal analysis. We then quantify the effects of key elements in the generation process, including decoding strategy, model size, model architecture, and prompt formulation, on the likelihood of smell occurrence. These causal analyses identify the factors that meaningfully influence smell propensity and guide the design of mitigation strategies that can be applied during inference. Finally, we complement this analysis with a user study that evaluates whether PSC helps developers interpret and assess generated code.
@@
We provide a causal analysis quantifying the influence of generation strategy, model size, model architecture, and prompt formulation.
@@
We release a dataset, benchmark, and implementation to support future research on explaining and mitigating smell generation (psc_repo).
@@ 4. Methodology
[Usefulness] How do developers assess code smells in LLM-generated code when supported by PSC? We conduct a user study to examine whether PSC helps developers evaluate generated code, focusing on its perceived relevance, its impact on confidence, and its usefulness in identifying introduced smells.
@@ 4.1. Semantic Preserving Code Transformations
Our methodology draws on prior work in robustness evaluation for neural program repair using natural transformations (sept_paper). We focus specifically on statement-level transformations. Broader edits, such as converting while loops to for loops (While2For), can significantly alter the code’s structural context, thereby influencing the detection of code smells. Limiting the analysis to statement-level changes preserves semantics and overall structure while still introducing meaningful syntactic variation.